# On Listwise Reranking for Corpus Feedback

### v1

@@ Abstract.
$c$s: Information systems Information retrieval
Figure 1. A cost-performance trade-off of rerankers (nDCG@10; see Appendix Table 4 for full results).
@@ 1. Introduction
Reranking captures document interactions as a form of corpus feedback to improve retrieval performance (Mao et al., 2024; Pradeep et al., 2023; Rathee et al., 2025a). To illustrate, listwise rerankers reflect local interactions by reranking documents within a window. For a large input, this window shifts forward to compare high-ranking documents with the next unseen documents, until the window reaches the other end of list. However, such baselines, due to serialized window shifts, fails to capture interaction feedback across non-adjacent windows.
One extreme of enriching corpus feedback is SlideGAR (Rathee et al., 2025a), which assumes a graph capturing all pairwise document interactions then leverages the neighborhood information beyond the window from the graphs. While effective, this approach is limited by the availability of such graphs, requires full corpus access to construct them, and incurs substantial cost from repeated bi-encoder calls (e.g., TCT (Lin et al., 2021b)), as well as a prohibitive $\mathcal{O}(N^{2})$ memory footprint for a corpus of size $N$.
In another extreme, graph-free rerankers such as TourRank (Chen et al., 2025a) conduct multiple rounds of tournaments across windows. Despite delivering a strong performance with SlideGAR without requiring explicit graphs, TourRank incurs heavy computation due to the large number of LLM calls required to remain competitive.
Figure 1 summarizes this tradeoff landscape by plotting average effectiveness against LLM calls. SlideGAR lies in the upper‑left: strong performance with sliding‑window baseline budget (9 calls), but requires additional compute and memory to construct the corpus graph.
Another extreme in upper-right is tournament‑style listwise reranker (e.g., TourRank) that is graph-free, improves performance but requires far more calls (about 127).
@@ 2. Related Works
Graph-based reranker Graph-aware methods such as SlideGAR (Rathee et al., 2025a) exploit explicit doc-doc relations to guide adaptive reranking, achieving strong quality, but requiring pre-built corpus graphs with notable memory and pre-processing costs. Our work keeps the call budget low without requiring a doc-doc retriever by reconstructing a sparse interaction graph directly from past listwise rankings and reusing it in SlideGAR.
Graph-free reranker However, a corpus or graph is not always available. Then, advanced reranking for ranked results can be used to overcome the gap, albeit at an additional cost (Rathee et al., 2025b). We focus on listwise rerankers (Pradeep et al., 2023) that jointly reason over candidate sets. They often rely on sliding windows or tournament-style comparisons. RankZephyr (Pradeep et al., 2023) is a strong fine-tuned listwise ranker; ReaRank (Zhang et al., 2025) emphasizes reasoning ability with more inference time; TourRank (Chen et al., 2025a) improves effectiveness at the cost of more LLM calls. Contemporary advanced ranking approaches (Liu et al., 2025, 2024b, 2024a) optimize in ICL/demo selection, collaborate small/large rankers with order adjustment, or expand context windows. Our distinction is that we frame listwise methods as implicitly inducing local interaction graphs among candidates with zero LLM-call overhead.
Query-stream reuse Transforming an entire corpus into a graph is a costly process, which (Chen et al., 2025b) argued is unnecessary. Our work shares this motivation, leveraging batch and stream processing (Ding et al., 2011; Eslami et al., 2019, 2020; Mackenzie and Moffat, 2023) to demonstrate gains from shared computation and intermediate results across queries.. Our incremental L2G updates operationalize this principle for reranking: listwise signals accumulate into a sparse graph that stays effective under order perturbations (RQ3).
@@ 3. Method: Listwise signals to Graph (L2G)
We study treating reranker outputs as signals and aggregate document pairs and their co-occurrence with queries, which induce a first-order similarity matrix that serves as a proxy for an implicit graph over documents.
3.1. Setup
Fig. 2 visualizes the building process. Let $\mathcal{Q}=\{q_{1},\dots,q_{m}\}$ be a stream of user queries. For each $q_{i}$ a listwise reranker receives a candidate set $\mathcal{C}_{i}=\{d_{i,1},\dots,d_{i,k}\}$ and returns a total order $\pi_{i}:\mathcal{C}_{i}\!\to\![k]$, where $\pi_{i}(d)$ is the rank of document $d$ for query $q_{i}$ (one-dimensional). We denote by $\mathcal{D}=\bigcup_{i}\mathcal{C}_{i}$ the (ever‑growing) global document pool.
@@ 3.2. From Rankings to a First-order Graph
First, we introduce a method for converting each ranked list for $q_{i}$ into a score vector. When accumulated, this forms a vectorized matrix for multiple queries, which becomes the basis for the underlying doc-doc graph structures. We map each ranked list into a dense document score vector $\mathbf{a}_{i}\in\mathbb{R}^{|\mathcal{D}|}$ with $\mathbf{a}_{i}]_{d}=(k-\pi_{i}(d)+1)$ if $d\in\mathcal{C}_{i}$, and 0 otherwise. In other words, the scores get assigned from highest to lowest depending on their rank for each query. We approximate pairwise document affinity as $\mathbf{D}^{(1)}=\mathbf{A}\mathbf{A}^{\top}$, where $D^{(1)}_{d_{1},d_{2}}=\sum_{i}[\mathbf{a}_{i}]_{d_{1}}\,[\mathbf{a}_{i}]_{d_{2}}$ accumulates evidence over all queries in which both documents co-occur, and treat this $\mathbf{D}^{(1)}$ as an approximation of the pairwise document similarity matrix. Related works suggest that greater query diversity leads to more accurate approximations (Kumar et al., 2012; Uhrenholt et al., 2021).
@@ 3.3. Handling Sparsity and Higher‐Order Connectivity
The first-order similarity $\mathbf{D}^{(1)}$ is often sparse due to the limited overlap between candidate sets across queries. However, multi-hop connections—e.g., two documents may not co-occur but may each co-occur with a shared third document—can capture higher-order similarity structure. Formally, we extend the first-order graph by propagating affinity via $k$-hop random walks: $\mathbf{D}^{(k)}=\mathbf{D}\mathbf{D}\cdots\mathbf{D}$ ($k$ times), analogous to power iteration in PageRank (Page et al., 1999). Larger $k$ increases recall but risks rank collapse towards a uniform stationary distribution. We therefore (i) renormalize rows to unit $k_{1}$ norm after each multiplication and (ii) cap $k\leq 3$ in practice.
@@ 3.5. Online Updates for Unseen Documents
The challenge in maintaining the document matrix is that in practice, the document pool grows dynamically as new queries introduce previously unseen documents, making it impractical to recompute the entire similarity matrix $\mathbf{D}^{(1)}$ from scratch each time. To support pairwise similarity for uncovered or newly added documents, As a solution, we develop an incremental update strategy that efficiently extends the graph structure without full recomputation.
Formally, the document pool approximates the document universe unknown a priori. As a dynamic approximation, when a new batch of queries adds a set $\Delta\mathcal{D}$ of unseen documents, we update $\mathbf{A}_{\text{new}}=\begin{bmatrix}\mathbf{A}\\ \mathbf{A}_{\Delta}\end{bmatrix}$ and maintain $\mathbf{D}^{(1)}$ in block form instead of re-computing from scratch: $\mathbf{D}^{(1)}_{\text{new}}=\begin{bmatrix}\mathbf{D}^{(1)}&\mathbf{B}\\ \mathbf{B}^{\!\top}&\mathbf{C}\end{bmatrix}$, where $\mathbf{B}=\mathbf{A}\mathbf{A}_{\Delta}^{\top}$ and $\mathbf{C}=\mathbf{A}_{\Delta}\mathbf{A}_{\Delta}^{\top}$. This costs $\mathcal{O}(|\mathcal{D}|\,|\Delta\mathcal{D}|\,m)$ rather than $\mathcal{O}(|\mathcal{D}_{\text{new}}|^{2}m)$. Since $\Delta|\mathcal{D}|\ll|\mathcal{D}|$ in typical streaming scenarios, this provides substantial computational savings.
@@ 3.6. Complexity and Memory Footprint
A key advantage of L2G is its elimination of expensive doc-doc similarity computations that plague existing graph-based methods. Unlike approaches such as SlideGAR-TCT that require explicit bi-encoder calls for every document pair, L2G introduces no additional LM calls for doc-doc similarity and downstream modules simply read the needed relations directly from the on-the-fly graph G.
Operationally, the method requires expansion of $D$ upon the arrival of a new ranked list, and applies three hops of propagation to form $\mathbf{D}^{(3)}$ when a new query’s ranking arrives. In terms of space, L2G stores only the sparse graph G (plus minimal ID maps), never materializing doc–doc matrices or document embeddings while entirely avoiding the need for any GPU-dependent doc-doc retriever. Direct comparison of L2G over SlideGAR--TCT under a matched setup are further discussed at Sec. 4.3.
@@ 4. Results
Our empirical study addresses three overarching questions. First, we ask about effectiveness: can L2G, built only from past listwise rankings, recover doc-doc structure closely enough to match graph-aware oracles such as doc-doc affinity and full-corpus SlideGAR? Second, we examine efficiency and cost: does L2G keep the number of LLM calls fixed at the sliding-window budget while offering favorable latency and memory scaling compared to graph-based baselines? Finally, we assess robustness and generality: is L2G stable under variations in query order and coverage, and does it transfer across different rerankers and first-stage retrievers?
@@ 4.1. Experimental Setup
Datasets & Evaluation. We evaluate on MS MARCO TREC (Craswell et al., 2020) DL’19/’20 (43/54 queries) and a subset of BEIR (Thakur et al., 2021) collections (TREC-COVID, Signal, TREC-NEWS, and Touche) with query size less than 100, and report nDCG@10 (Järvelin and Kekäläinen, 2002) under reranking budget $c$ (c= 100 or 1000) rounded to the nearest tenth, mostly prepared with the pre-built Pyserini index (Lin et al., 2021a).
Retrieval & Ranking Models. We use BM25 (Robertson et al., 2009) and extend to Contriever (Izacard et al., 2022) as the first-stage retriever. Listwise rerankers include fine-tuned RankZephyr (Pradeep et al., 2023) and a reasoning model ReaRank (Zhang et al., 2025), operating with sliding windows (size=20, step=10) that output permutations without explicit scores.
Baselines. We compare to: (i) non-adaptive Sliding Window, and 3 different (ii) Graph-based SlideGAR--TCT baselines. (1) (full-corpus) assumes oracle access to a pre-computed TCT (Lin et al., 2020) graph over the entire collection. (2) (doc–doc affinity) is a more practical baseline, which mimics the online graph construction scenario just like ours (Sec. 3.3), limiting the graph neighbors to a local graph built only from the query’s retrieved top-$c$ candidates, e.g., top-100. Lastly, (iii) (Random) randomizes neighbor ordering from local top-$c$ candidates, verifying that L2G’s gains are not due to chance.
Compute. All experiments were executed on either a workstation equipped with eight 24 GB RTX 3090 GPUs, or eight 48 GB NVIDIA A6000 GPUs.
@@ 6. Ethical Considerations
While the L2G is far more efficient than SlideGAR, additional requirements for graph construction may still limit accessibility for resource-constrained organizations. The efficiency of L2G can be further optimized by tuning graph update intervals at the cost of performance, or through implementation of the GPU version of L2G. Our simple implementation leaves additional speed/memory gains as future work, which could help address these accessibility concerns.
Appendix A Appendix
Doc Ranker D19 D20 D21 D22 D23 D-H AVG cov. news nfc sig. r04 touc. dbp SciF. BEIR ALL #Calls
@@ Appendix A Appendix
Tab 4 refers to the full results on the comparison between advanced listwise reranking methods v.s. slidegar (full corpus).

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USAbooktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAdoi: 10.1145/3773966.3779404isbn: 979-8-4007-2292-9/2026/02ccs: Information systems Information retrievalccs: Information systems Retrieval models and rankingccs: Information systems Language modelsccs: Information systems Learning to rank
Figure 1. A cost-performance trade-off of rerankers (nDCG@10; see Table 4 for full results).
@@ 1. Introduction
Reranking captures document interactions as a form of corpus feedback to improve retrieval performance (mao-etal-2024-rafe; rankzephyr; slidegar). To illustrate, listwise rerankers reflect local interactions by reranking documents within a window. For a large input, this window shifts forward to compare high-ranking documents with the next unseen documents, until the window reaches the other end of the list. However, such baselines, due to serialized window shifts, fail to capture interaction feedback across non-adjacent windows.
One extreme of enriching corpus feedback is SlideGAR (slidegar), which assumes a graph capturing all pairwise document interactions, then leverages the neighborhood information beyond the window from the graphs. While effective, this approach requires full corpus graph access to construct them, and incurs substantial cost from repeated bi-encoder calls (e.g., TCT (lin-etal-2021-tctcolbert)), as well as a prohibitive $\mathcal{O}(N^{2})$ memory footprint for a corpus of size $N$. In another extreme, graph-free rerankers such as TourRank (tourrank) conduct multiple rounds of tournaments across windows. Despite delivering a strong performance with SlideGAR without requiring explicit graphs, TourRank incurs heavy computation due to the large number of LLM calls required to remain competitive.
Figure 1 summarizes this tradeoff landscape by plotting average effectiveness against LLM calls. SlideGAR lies in the upper‑left: strong performance with sliding‑window baseline budget (9 calls), but requires additional compute and memory to construct the corpus graph. Another extreme in the upper-right is tournament‑style listwise reranker (e.g., TourRank) that is graph-free, improves performance but requires far more calls (about 127).
@@ 2. Related Works
Graph-based reranker Graph-aware methods such as SlideGAR (slidegar) exploit explicit doc-doc relations to guide adaptive reranking. While they achieve strong quality, they requiring pre-built corpus graphs with huge memory and pre-processing costs. Our work keeps the call budget low without requiring a doc-doc retriever by reconstructing a sparse interaction graph directly from past listwise rankings and reusing it in SlideGAR.
Graph-free reranker However, a corpus or graph is not always available. Then, advanced reranking for ranked results can be used to overcome the gap, albeit at an additional cost (rathee2025testtimecorpusfeedbackretrieval). We focus on listwise rerankers (rankzephyr) that jointly reason over candidate sets. They often rely on sliding windows or tournament-style comparisons. RankZephyr (rankzephyr) is a strong fine-tuned listwise ranker; ReaRank (rearank) emphasizes reasoning ability with more inference time; TourRank (tourrank) improves effectiveness at the cost of more LLM calls. Contemporary advanced ranking approaches (coranking; demorank; slidingwindowsnotend) optimize in ICL/demo selection, collaborate small/large rankers with order adjustment, or expand context windows. Our distinction is that we frame listwise methods as implicitly inducing local interaction graphs among candidates with zero LLM-call overhead.
Query-stream reuse Transforming an entire corpus into a graph is a costly process, while previous work  (chen2025dontneedprebuiltgraphs) argued is unnecessary. Our work shares this motivation, leveraging batch and stream processing (Ding2011batchqueryprocessing; eslami2019psidb; eslami2020query; mackenzie2023index) to demonstrate gains from shared computation and intermediate results across queries. Our incremental L2G updates operationalize this principle for reranking: listwise signals accumulate into a sparse graph that stays effective under order perturbations.
@@ 3. Method: Listwise signals to Graph (L2G)
We study treating reranker outputs as document relationship signals and aggregate document pairs and their co-occurrence with queries, which induces a first-order similarity matrix that serves as a proxy for an implicit graph over documents. Fig. 2 visualizes the building process. Let $\mathcal{Q}=\{q_{1},\dots,q_{m}\}$ be a stream of user queries. For each $q_{i}$ a listwise reranker receives a candidate set $\mathcal{C}_{i}=\{d_{i,1},\dots,d_{i,k}\}$ and returns a total order $\pi_{i}:\mathcal{C}_{i}\!\to\![k]$, where $\pi_{i}(d)$ is the rank of document $d$ for query $q_{i}$ (one-dimensional). We denote by $\mathcal{D}=\bigcup_{i}\mathcal{C}_{i}$ the (ever‑growing) global document pool.
@@ 3.1. From Rankings to a First-order Graph
First, we introduce a method for converting each ranked list for $q_{i}$ into a score vector. When accumulated, this forms a vectorized matrix for multiple queries, which becomes the basis for the underlying doc-doc graph structures. We map each ranked list into a dense document score vector $\mathbf{a}_{i}\in\mathbb{R}^{|\mathcal{D}|}$ with $\mathbf{a}_{i}]_{d}=(k-\pi_{i}(d)+1)$ if $d\in\mathcal{C}_{i}$, and 0 otherwise. In other words, the scores get assigned from highest to lowest depending on their rank for each query. We approximate pairwise document affinity as $\mathbf{D}^{(1)}=\mathbf{A}\mathbf{A}^{\top}$, where $D^{(1)}_{d_{1},d_{2}}=\sum_{i}[\mathbf{a}_{i}]_{d_{1}}\,[\mathbf{a}_{i}]_{d_{2}}$ accumulates evidence over all queries in which both documents co-occur, and treat this $\mathbf{D}^{(1)}$ as an approximation of the pairwise document similarity matrix. Related works suggest that greater query diversity leads to more accurate approximations (nystrom; inducingpoints).
@@ 3.2. Handling Sparsity and Higher‐Order Connectivity
The first-order similarity $\mathbf{D}^{(1)}$ is often sparse due to the limited overlap between candidate sets across queries. However, multi-hop connections—e.g., two documents may not co-occur but may each co-occur with a shared third document—can capture higher-order similarity structure. Formally, we extend the first-order graph by propagating affinity via $k$-hop random walks: $\mathbf{D}^{(k)}=\mathbf{D}\mathbf{D}\cdots\mathbf{D}$ ($k$ times), analogous to power iteration in PageRank (PageRank1999ThePC). Larger $k$ increases recall but risks rank collapse towards a uniform stationary distribution. We therefore (i) renormalize rows to unit $k_{1}$ norm after each multiplication and (ii) cap $k\leq 3$ in practice.
@@ 3.4. Online Updates for Unseen Documents
The challenge in maintaining the document matrix is that in practice, the document pool grows dynamically as new queries introduce previously unseen documents, making it impractical to recompute the entire similarity matrix $\mathbf{D}^{(1)}$ from scratch each time. To support pairwise similarity for uncovered or newly added documents, As a solution, we develop an incremental update strategy that efficiently extends the graph structure without full recomputation. Formally, the document pool approximates the document universe unknown a priori. As a dynamic approximation, when a new batch of queries adds a set $\Delta\mathcal{D}$ of unseen documents, we update $\mathbf{A}_{\text{new}}=\begin{bmatrix}\mathbf{A}\\ \mathbf{A}_{\Delta}\end{bmatrix}$ and maintain $\mathbf{D}^{(1)}$ in block form instead of re-computing from scratch: $\mathbf{D}^{(1)}_{\text{new}}=\begin{bmatrix}\mathbf{D}^{(1)}&\mathbf{B}\\ \mathbf{B}^{\!\top}&\mathbf{C}\end{bmatrix}$, where $\mathbf{B}=\mathbf{A}\mathbf{A}_{\Delta}^{\top}$ and $\mathbf{C}=\mathbf{A}_{\Delta}\mathbf{A}_{\Delta}^{\top}$. This costs $\mathcal{O}(|\mathcal{D}|\,|\Delta\mathcal{D}|)$ rather than $\mathcal{O}(|\mathcal{D}_{\text{new}}|^{2})$. Since $\Delta|\mathcal{D}|\ll|\mathcal{D}|$ in typical streaming scenarios, this provides substantial computational savings.
@@ 3.5. Complexity and Memory Footprint
A key advantage of L2G is its elimination of expensive doc-doc similarity computations that plague existing graph-based methods. Unlike approaches such as SlideGAR-TCT that require explicit bi-encoder calls for every document pair, L2G introduces no additional LLM calls for doc-doc similarity and downstream modules simply read the needed relations directly from the on-the-fly graph G.
Operationally, the method requires expansion of $D$ upon the arrival of a new ranked list, and applies three hops of propagation to form $\mathbf{D}^{(3)}$ when a new query’s ranking arrives. In terms of space, L2G stores only the sparse graph G (plus minimal ID maps), never materializing doc–doc matrices or document embeddings while entirely avoiding the need for any GPU-dependent doc-doc retriever. Direct comparison of L2G over SlideGAR--TCT under a matched setup is further discussed at Sec. 4.3.
@@ 4. Results
Our empirical study addresses three questions. 1. Effectiveness: can L2G, built only from past listwise rankings, recover doc-doc structure closely enough to match graph-aware oracles such as doc-doc affinity and full-corpus SlideGAR? 2. Efficiency and cost: does L2G keep the number of LLM calls fixed at the sliding-window budget while offering favorable latency and memory scaling compared to graph-based baselines? 3. Robustness and generality: is L2G stable under variations in query order and coverage, and does it transfer across different rerankers and first-stage retrievers?
@@ 4.1. Experimental Setup
Datasets & Evaluation. We evaluate on MS MARCO TREC (craswell2020overview) DL’19/’20 (43/54 queries) and a subset of BEIR (beir) collections (TREC-COVID, Signal, TREC-NEWS, and Touche) with query size less than 100, and report nDCG@10 (ndcg) under reranking budget $c$ (c= 100 or 1000) rounded to the nearest tenth, mostly prepared with the pre-built Pyserini index (pyserini).
Retrieval & Ranking Models. We use BM25 (bm25) and extend to Contriever (contriever) as the first-stage retriever. Listwise rerankers include RankZephyr (rankzephyr) andtextitReaRank (rearank), operating with sliding windows (size=20, step=10) that output permutations without explicit scores.
Baselines. We compare to: (i) non-adaptive Sliding Window, and 3 different (ii) Graph-based SlideGAR–TCT baselines. (1) full-corpus assumes oracle access to a pre-computed TCT (tct) graph over the entire collection. (2) doc–doc affinity is a more practical baseline, which mimics the online graph construction scenario just like ours (Sec. 3.2), limiting the graph neighbors to a local graph built only from the query’s retrieved top-$c$ candidates, e.g., top-100. Lastly, (iii) Random randomizes neighbor ordering from local top-$c$ candidates, verifying that L2G’s gains are not due to chance.
@@ 5. Conclusion
Acknowledgements.
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], and the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2024-00414981),
@@ 6. Ethical Considerations
While the L2G is far more efficient than SlideGAR, additional requirements for graph construction may still limit accessibility for resource-constrained organizations. The efficiency of L2G can be further optimized by tuning graph update intervals at the cost of performance, or through implementation of the GPU version of L2G. Our simple implementation leaves additional speed and memory gains as future work, which could help address these accessibility concerns.
Appendix
Doc Ranker D19 D20 D21 D22 D23 D-H AVG covid news nfc sig. r04 touc. dbp SciF. BEIR ALL #Calls
@@
Table 4 refers to the full results on the comparison between advanced listwise reranking methods v.s. slidegar (full corpus).
Appendix A Additional Discussion on the Efficiency of L2G
Building on Sec. 4.3, we further discuss the efficiency of L2G, with a particular focus on how its session-induced graph scales compared to baselines that rely on a corpus-level document–document affinity graph (e.g., SlideGAR-TCT).
Setup and fairness.
To keep the comparison conservative and even slightly favorable to the document–document affinity baseline (SlideGAR-TCT), we (i) enable document caching and GPU batching for TCT, and (ii) use 768-dimensional FP32 vectors for both queries and documents across all systems. All measurements are obtained on a GPU server with dual Intel Xeon Gold 6254 CPUs (36 cores total; 72 threads), 607 GB system memory, and 8 NVIDIA GeForce RTX 3090 GPUs (24 GB each). For reproducibility, the code used to obtain the efficiency results is available in the following GitHub branch. [^2] https://github.com/soyoung97/l2g-submit/tree/efficiency
A.1. Compute with respect to document pool size.
Compute (wall time).
Since the graph grows dynamically as new queries add signals from previously unseen documents, the computational cost of the graph propagation step ($G\sim D^{3}$) and online updates does grow as the document pool scales. With k=20k{=}20 and eager graph updates (U=1U{=}1), L2G shows near-linear per-query latency in the number of processed queries $q$:
$\ell_{\mathrm{L2G}}(q)\approx 0.035\,q~\text{s}.$
However, we believe there is substantial room for improvement. This can be further optimized by using advanced caching techniques and efficient graph storage algorithms, such as pre-allocating document slots in batches to accommodate anticipated graph growth. Currently, the graph is updated whenever a single new ranking signals arrive; by batching or delaying graph updates (i.e., increasing the update interval $U$), the constant in the 0.035q0.035\,q slope would shrink proportionally, yielding faster runtime without changing the scoring model. We leave such optimizations as future work.
Memory.
Peak working memory for L2G also scales linearly with the number of processed queries $q$:
$\mu_{\mathrm{L2G}}(q)\approx 0.0276\,q~\text{MB}.$
In contrast, the full-corpus doc-doc affinity variant of SlideGAR (without Top-100 truncation) incurs memory size and graph construction latency that scale tremendously with corpus size. A key advantage of L2G is that its resource consumption depends only on $q$, not on the full corpus size.
Takeaways.
(i) L2G attains SlideGAR-level quality without any offline graph building; (ii) incremental sparse updates reduce compute and memory versus naïve $|\mathcal{D}|^{2}$ construction and provide a tunable cost-latency knob ($U$) that is largely orthogonal to model quality; (iii) even in a SlideGAR-favored configuration (document caching and GPU batching enabled), L2G is faster and lighter for small to medium sessions, and remains competitive thereafter. Overall, cost parity with SW and quality parity with doc-doc affinity make L2G an attractive default when storing a small sparse graph is acceptable.
Appendix B Limitations
Our method’s effectiveness is highly affected on documents co-occurring in retrieved lists, making it less effective in scenarios with low query-document overlap. In this case, when the constructed graph remains sparse, L2G may not provide substantial benefit over simpler baselines.